A glass Louvre pane 0.4 cm thick has an area 2×10^4 cm^2. On a very cold day the temperature difference between the inside and the outside surfaces of the pane is 25 degree Celsius determine the heat flow through the window
1
Expert's answer
2020-06-29T14:22:05-0400

As per the question,

Thickness of the pane=0.4cm

Area"=2\u00d710^4 cm^2"

Temperature difference between the pan "(\Delta T)=25^\circ C"

Let the conductivity of the material "=K"

Hence, rate of the heat flow "\frac{dQ}{dt}=\frac{KAd\theta}{l}"

Now, substituting the values,

"\Rightarrow \dfrac{dQ}{dt}=\frac{K\times 25\times2\times 10^6 cm^2}{0.4cm}=125\times K
\times 10^6 J/sec"